An iup runner has reduced her time for the Indiana-nowhere run from 5hr to4hr.She also has increased her speed by 2mph. The distance of the Indiana-nowhere run is
A 8 mi b 16mi c 20 mi d 32 mi e none of these

Answer by richwmiller(17219)

r*t=d
5x=4(x+2)
5x=4x+8
x=8 miles
check
5*8=40
4*10=40
ok